OpenStudy (anonymous):

what is -1/2x+2=-x+7

OpenStudy (kropot72):

\[-\frac{1}{2}x+2=-x+7\] What do you get when you add x to both sides?

OpenStudy (anonymous):

you get - 1/2

OpenStudy (kropot72):

When x is added to both sides you get \[x-\frac{1}{2}x+2=x-x+7\] Now you can simplify both sides